xx CarpetMask

CarpetMask

Erik Schnetter <schnetter@cct.lsu.edu>

April 20 2007

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11excludexexterior 10Scope: privatex01 BOOLEAN

11Description: exclude the exterior of this radius, otherwise the interior

11                  10  x01  Default: no

11

11excludedxcentrexx 10Scope: privatex01  REAL

11Description: x coordinate of excluded region

11  Range  10  x01  Default: 0.0

11*:*  10

11

11excludedxcentrexy 10Scope: privatex01  REAL

11Description: y coordinate of excluded region

11  Range  10  x01  Default: 0.0

11*:*  10

11

11excludedxcentrexz 10Scope: privatex01  REAL

11Description: z coordinate of excluded region

11  Range  10  x01  Default: 0.0

11*:*  10

11

11excludedxradius  10Scope: privatex01  REAL

11Description: radius of excluded region

11 Range 10 x01 Default: -1.0
11-1.0  10do not exlude this region
110.0:*  10

11

11excludedxsurface  10Scope: privatex01  INT

11Description: index of excluded surface

11 Range 10 x01 Default: -1
11-1  10do not exclude

110:*  10

11

11excludedxsurfacexfactor 10Scope: privatex01  REAL

11Description: shrink factor of excluded surface

11  Range  10  x01  Default: 1.0

110:*  10

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Produce screen output

11                  10  x01  Default: no

11

11maxnphi  10Scope: shared from SPHERICALSURFACEx01INT

11

11maxntheta  10Scope: shared from SPHERICALSURFACEx01INT

11

11nsurfaces  10Scope: shared from SPHERICALSURFACEx01INT

11

11symmetricxx  10Scope: shared from SPHERICALSURFACEx01BOOLEAN

11symmetricxy  10Scope: shared from SPHERICALSURFACEx01BOOLEAN

11

11symmetricxz  10Scope: shared from SPHERICALSURFACEx01BOOLEAN

11

7 Interfaces

General Implements: carpetmask

Inherits: grid sphericalsurface

Uses header: CarpetReducexbits.h

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetMask. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE Scheduled Functions

CCTKxPARAMCHECK x carpetsurfaceparamcheck x check parameters x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxPOSTSTEP x maskbasexsetupmask x set up the weight function x

11x11After:11sphericalsurfacexhasbeenset
11 x 11Type:11group

11

SetupIMask x carpetexcludedsetup x set up the weight function for the excluded regions x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Reads: 11grid::x
11 x 11x  11grid::y
11 x 11x  11grid::z
11 x 11Type: 11function

11

SetupIMask x carpetsurfacesetup x set up the weight function for the excluded spherical surfaces x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Reads: 11grid::x
11 x 11x  11grid::y
11 x 11x  11grid::z
11 x 11x  11sphericalsurface::sfxcoordinatexdescriptors
11 x 11x  11sphericalsurface::sfxshapexdescriptors
11 x 11x  11sphericalsurface::sfxinfo
11 x 11x  11sphericalsurface::sfxactive
11 x 11x  11sphericalsurface::sfxorigin
11 x 11x  11sphericalsurface::sfxradius
11 x 11Type: 11function